solve equation 2x+y=6 and x+2y=9 , x= ,y=
Answer by 1998Baby(63)

I'm going to use the elimination method for this
2x+y=6
x+2y=9+> multiply only the bottom row by -2
____
you should get
2x+y=6
-2x-4y=-18
you can cross out the x values and then add downward.
you get
-3y=-12
divide by -3 to -12
y=4
now take y=4 and put it in any equation
2x+(4)=6
-4 on both sides.
you get 2x=2
divide by 2
x=1
____
y=4
x=1
____
hope this helps :)